Nitrogen and hydrogen combine at a high temperature, in the presence of a catalyst, to produce ammonia.

N2(g)+3H2(g)⟶2NH3(g)

Assume 0.200 mol N2 and 0.647 mol H2 are present initially.

After complete reaction, how many moles of ammonia are produced?

NH3: mol

How many moles of H2 remain?

H2: mol

How many moles of N2 remain?

N2: mol
Answer #1

According to the reaction,

1 mole N2 require 3 mole H2

0.200 mole N2 require 3×0.200 = 0.600 mole H2

Since we have more than this hence N2 is limiting reagent.

1 mole N2 give 2 mole NH3

0.200 mole N2 give 0.200×2 = 0.400 mole NH3

Moles of N2 remain = 0 moles

Moles of H2 remain = 0.647-0.600 = 0.047 mole

Please give feedback!